Fermi National Laboratory
 
Operation Summary for January 30, 2012 - February 1, 2012
 

Monday January 30

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI and MiniBooNE experiments taking beam.

During the day shift, Operators and the Duty Mechanic responded to a report of a pump leak in the Linac basement at 8:34 AM; the pump was for the Linac hot water system and the Duty Mechanic brought the backup pump online. At 8:41 AM, a Linac RF station (LRF4) tripped off; a Linac technician brought it back online. A Booster RF station (BRF11) tripped off at 9:17 AM; it reset. Operations turned on and tuned up the Muon Rings at 9:56 AM for studies. At 10:08 AM, BRF16 tripped off; it reset. The Meson FTBF experiment T-1018 asked to be searched and secured at 10:09 AM, but FESS personnel were still in Enclosure C working on a sump pump; they came out and will finish their work later. Operations reported at 10:32 AM that they were having a problem turning on a MTest power supply (HP3DS); a Power Supply person investigated and reported it had a LCW leak; he began repairs. A Booster person began a study at 11:05 AM. At 12:04 PM, BRF1 tripped off; Booster RF personnel investigated and then brought it back online. The Power Supply person working on the HP3DS completed his repairs and Operations established beam to T-1018 at 12:18 PM. BRF1 tripped off at 12:55 PM; RF personnel investigated and brought the station back online. A Booster vacuum pump (B:IPRF22) tripped off and brought down BRF11 and 12 at 2:26 PM. At 3:30 PM, BRF3, 4, and 14 all tripped off on a RF gap fault indication; they all reset. Operators believed that the trips were due to the Booster studies.

During the evening shift, at 4:05 PM, a Linac quadrupole magnet power supply (L:QPS104) tripped off; it reset. The Muon Rings studies ended at 4:38 PM. The Booster studies ended at 5:01 PM. At 6:34 PM. LRF5 tripped off; it reset. Operators reported at 7:48 PM that they had completed a 750 KeV line tune. Operations completed a Booster injection tune at 10:03 PM. Operations disabled beam to T-1018 at 10:05 PM; this ended the experiment’s current run at Fermilab. LRF2 tripped off at 11:30 PM; it reset.

Tuesday January 31

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE.

During the midnight shift, Operations halted all beam, except for Linac, at 2 AM in preparation for the seven-day shutdown to replace a MI quadrupole magnet. Operators had MI prepared for access by 2:47 AM. Operators had the Transfer Hall, F-Sector, and MiniBooNE prepared for access by 3:06 AM. Operators had the Minos Absorber and Alcoves prepared for access by 3:23 AM. Operators had Booster prepared for access by 5:35 AM. Operators began handing out keys for the MI access at 6 AM. At 7:29 AM, Neutron Therapy Facility (NTF) personnel began preparing to treat one patient later in the morning.

During the day shift, Operators had M01, M02, M03, M04, and M05 ready for access by 9:40 AM. NTF personnel reported at 10:30 AM that they had completed treatment on their only patient for the day. MI personnel reported at 12:49 PM that they had disconnected the Recycler quadrupole magnets (R:QT701, 702, 703, and 704) from their power supply inside the MI-12 service building so the magnets could be safely removed.

During the evening shift, a Mechanical Technician reported finding water under a power supply at MI-40; Operators will pass this on power supply techs on Wednesday morning. Operators reported at 11:34 PM that Welders had completed their welding on the downstream end of Q403’s beam pipe in MI.

Wednesday February 1

The midnight shift began with the accelerator complex in shutdown for the MI quadrupole magnet replacement.

During the midnight shift, at 3:11 AM, LRF4 tripped off; it reset. Operators also reset LRF3.

Operations Coordinator (OpsCo) Plans for Today

The OpsCo reports that Controls personnel will be replacing hardware for the Time Line Generator hardware this morning. Linac will also be conducting maintenance during the shutdown, but they will come up for NTF treatments.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/01/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory